The density of gold is 19 times that of water. If you take a gold crown weighing 20 N and submerge it in water, how much upward force must you exert on the submerged crown to keep it from accelerating?
Answer #1

To keep it from accelerating, net force must be zero

downward force = 20N

upward buoyancy force = 20 * density of water / density of gold = 20 * 1/19 = 20/19 N

so upward force need to be applied = 20 - 20/19 = 20(1-1/19) = 360/19 N or 18.95 N